A basketball player scores 23 times during one game. She scored a total of 37 points, two for each two-point shot and one for each free throw, How many two-point shots did she make> How many free throws?
Simplify the answers

Answer by checkley79(3341)

A+B=23
A=23-B
2A+1B=37
2(23-B)+1B=37
46-2B+1B=37
-1B=37-46
-B=-9
B=9 NUMBER OF ONE POINTERS.
A+9=23
A=23-9
A=14 NUMBER OF TWO POINTERS.
PROOF:
2*14+1*9=37
28+9=37
37=37